OpenStudy (anonymous):

What are two ways to simplify the square root of 320.? You can show a square root by writing "sqrt()". For example, the square root of 2 can be written as "sqrt(2)"

OpenStudy (mrnood):

you need to find factors of 320 that include 'square' numbers e.g. 4 * 80 sqrt (320) = sqrt(4 * 80) = sqrt(4) * sqrt (80) since 4 is 2*2 you can simplify that. There are bigger factors which work the same way